4. Suppose you can hire a mechanic to work on your car for up to six hours. The total benefit of repair work is B(H)-100H-5H2 and the total cost is C(H) = 10H + 5H2, where H is the number of hours. (a) What is the net benefit? (b) Graph the net benefit function from 0 to 10 hours (c) What is your marginal benefit? (d) What is your marginal cost? (e) Graph marginal benefit and marginal cost curves from 0 to 10 hours on the same graph (f) What is your choice that maximizes the net benefit? (g) What is the total benefit at the optimal choice? (h) What is the total cost at the optimal choice? (i) What is the net benefit at the optimal choice?
Answer #1

The total benefit is B(H) = 100H -5H and total cost is C(H)=10H +5H? Here. H is the number of hours. Find the net benefit. NeDerivate the benefit function with respect to H. MB(H) = ABCH) = 100-107] Thus, marginal benefit is 100-101 (d) Derivate theANB(H) -90-20H = 0 ΔΗ 20H = 90 H = 20 = 14,5 Thus, 4.5 hours where the net benefit maximizes. Find the total benefit at 4.5 hC(H)=10(4.5)+5(4.5)2 = 45 +101.25 = 146.25 Net benefit is the difference between the total benefit and total cost Net benefit